Can someone help me with this problem?
I need to find the GCF (greatest common factor) for each groups of numbers:
36, 54, and 180
Thanks,
Sher

Answer by stanbon(57967)

36, 54, and 180
Write each in prime-factor form, as follows:
36= 2^2 * 3^2
54= 2 * 3^3
180 = 2^2 * 3^2 * 5
The GCF must contain the lowest prime power factors common to all the numbers:
GCF = 2 * 3^2 = 18
It doesn't have a 2^2 because 2^2 is not common to all three numbers.
It doesn't have a 3^3 for the same reason.
It doesn't have a 5 for the same reason.
Cheers,
Stan H.